OpenStudy (blondisaurus):

Which of the following is a simplified expression equal to (comment section), for all x< -3

OpenStudy (blondisaurus):

\[\frac{ 9-x ^{2} }{ x-3 }\]

OpenStudy (anonymous):

You can factor the numerator as a difference of squares then cancel 9 - x² = (3 - x)(3 + x)

OpenStudy (anonymous):

and then (3 - x) is the same as -(x - 3)

OpenStudy (anonymous):

\[\frac{ (3-x) (3+x)}{ (x-3) }=\frac{ -(x-3)(3+x) }{ (x-3) }\]

OpenStudy (anonymous):

=-(3+x)=-3-x